OpenStudy (theedwardsfamily):

Please help. What is the slope of the line passing through the points (0, −5) and (4, 2)?

OpenStudy (anonymous):

7/4

OpenStudy (alexandervonhumboldt2):

use slope formula: \[\frac{ y_1-y_2 }{ x_1-x_2 }\] substitute the values and solve

OpenStudy (alexandervonhumboldt2):

@TheEdwardsFamily

OpenStudy (theedwardsfamily):

Yea, I got 1/4. @AlexandervonHumboldt2

OpenStudy (theedwardsfamily):

Am I right?

OpenStudy (anonymous):

it's actually y2 - y1 divided by x2 - x1. y2= 2 y1= -5 x2= 4 x1= 0

OpenStudy (alexandervonhumboldt2):

y2-y1/x1-x2 is same as y1-y2/x1-x2

OpenStudy (alexandervonhumboldt2):

it is -7/-4=7/4

OpenStudy (theedwardsfamily):

oh I see!

OpenStudy (alexandervonhumboldt2):

as you get : (-5-2)/(0-4)=-7/-4=7/4

OpenStudy (theedwardsfamily):

thnx @AlexandervonHumboldt2 !!!

OpenStudy (alexandervonhumboldt2):

np:) @kaylie.sullivan there are same: (y2-y1)/(x2-x1)=-(y2-y1)/-(x2-x1)=(-y2+y1)/(-x2+x1)=(y1-y2)/(x1-x2) hahaha same fomulas